Rule:×4+2
Jacob is making patterns using marbles such as the one below. Show by means of a flow chart and a table, how many marbles he will use for the 6th pattern, the 11th pattern, the 41st pattern and the 100th pattern. Write down the rule for this pattern.
1
Expert's answer
2020-07-01T18:22:11-0400

Rule:×4+2 implies nth term is given by : "a_n= 4n+2"

So , the pattern sequence goes like : 6,10,14,18,22,...

Thus, 6th pattern is 4*6+2= 26

11th pattern is 4*11+2= 46

41st pattern = 4*41+2=166

100th pattern =4*100+2 =402